# Ophiuchus: Scalable Modeling of Protein Structures through Hierarchical Coarse-graining SO(3)-Equivariant Autoencoders

### v1

@@ 3.2 Self-Interaction
Our Self-Interaction transform draws inspiration from MACE [Batatia et al. (2022)]. This transformation updates the feature vectors $\mathbf{V}^{0:{l_{\max}}}_{i}$ localized at the same anchoring node $i$. Importantly, it blends feature vectors $\mathbf{V}^{l}$ of varying orders $l$ by employing tensor products of the features with themselves. We offer two implementations of these tensor products to cater to different computational needs. For a comprehensive explanation, please refer to Appendix B.
@@
$\tilde{\mathbf{V}}^{0:{l_{\max}}}\leftarrow\mathbf{V}^{0:{l_{\max}}}$  normal-$\triangleright$ Save For Later
$\mathbf{V}^{0:{l_{\max}}}\leftarrow\mathbf{V}^{0:{l_{\max}}}\oplus\left(\mathbf{V}^{0:{l_{\max}}}\right)^{\otimes 2}$  normal-$\triangleright$ Tensor Product with itself
$\mathbf{V}^{0:{l_{\max}}}\leftarrow\mathrm{Linear}(\mathbf{V}^{0:{l_{\max}}})$ $\mathbf{V}^{0:{l_{\max}}}\leftarrow\textrm{MLP}(\mathbf{V}^{l=0})\cdot\mathbf{V}^{0:{l_{\max}}}$  normal-$\triangleright$ Scalar-Irrep Multiplication
$\mathbf{V}^{0:{l_{\max}}}\leftarrow\mathbf{V}^{0:{l_{\max}}}+\tilde{\mathbf{V}}^{0:{l_{\max}}}$  normal-$\triangleright$ If the Multiplicities Match
$\mathbf{V}^{0:{l_{\max}}}\leftarrow\textrm{Norm}(\mathbf{V}^{0:{l_{\max}}})$  normal-$\triangleright$ Normalize $l$-Wise
@@ 3.3 Sequence Convolution
We introduce a one-dimensional roto-translational equivariant convolutional kernel designed to operate on representations of SO(3) and $\mathbb{R}^{3}$. Given a kernel window size $K$ and stride $S$, we concatenate representations $\mathbf{V}^{0:{l_{\max}}}_{i-\frac{K}{2}:i+\frac{K}{2}}$ with the same $l$ value. Additionally, we include normalized relative vectors between anchoring positions $\mathbf{P}_{i-\frac{K}{2}:i+\frac{K}{2}}$. Following conventional CNN architectures, this concatenated representation undergoes a linear transformation. The scalars in the resulting representation are then converted into weights, which are used to combine window coordinates into a new coordinate. To ensure translational equivariance, these weights are constrained to sum to one.
@@
$\mathbf{P}\leftarrow\sum_{k=1}^{K}w_{k}\mathbf{P}_{k}$ $\tilde{\mathbf{P}}_{1:K,1:K}\leftarrow\bigoplus_{i,j}\frac{\mathbf{P}_{i}-\mathbf{P}_{j}}{\|\mathbf{P}_{i}-\mathbf{P}_{j}\|}$   normal-$\triangleright$ All relative vectors
$\mathbf{V}^{0:{l_{\max}}}\leftarrow\textrm{Linear}\Big{(}\mathbf{V}^{0:{l_{\max}}}_{1:K}\oplus\tilde{\mathbf{P}}_{1:K,1:K}\Big{)}$ return $(\mathbf{P},\mathbf{V}^{0:{l_{\max}}})$
@@
When $S>1$, sequence convolutions reduce the dimensionality along the sequence axis, yielding mixed representations and coarse coordinates. To reverse this process, we introduce a transpose convolution algorithm that geometrically reconstructs both the sequence and its positional coordinates. For further details, please refer to Appendix C.
@@ 3.4 Spatial Convolution
Spatial convolutions update representations through message passing within k-nearest spatial neighbors. Message representations incorporate SO(3) signals from the vector difference between neighbor coordinates, and we aggregate messages with permutation-invariant means. After aggregation, we linearly transform the vector representations into a single vector that is used to update the coordinates.
@@
$\hat{\mathbf{V}}^{0:{l_{\max}}}\leftarrow\mathbf{V}^{0:{l_{\max}}}_{i}$  normal-$\triangleright$ Save for Later
$(\tilde{\mathbf{P}},\tilde{\mathbf{V}}^{0:{l_{\max}}})_{1:k}\leftarrow\textrm{$k$-Nearest-Neighbors}(\mathbf{P}_{i},\mathbf{P}_{1:N})$ $(\phi,R,u)_{1:k}\leftarrow\textrm{Embeddings}(\tilde{\mathbf{X}}_{1:k}-\mathbf{P}_{i})$  normal-$\triangleright$ Angular/Radial Embeddings
$\mathbf{V}^{0:{l_{\max}}}\leftarrow\textrm{Linear}\left(\left(\sum_{k}u_{k}\textrm{MLP}(R_{k})\tilde{\mathbf{V}}^{0:{l_{\max}}}_{k}\otimes\phi_{k}\right)/\left(\sum_{k}u_{k}\right)\right)$ $\mathbf{V}^{0:{l_{\max}}}\leftarrow\textrm{Norm}\left(\mathbf{V}^{0:{l_{\max}}}\right)$  normal-$\triangleright$ Normalize $l$-Wise
$\mathbf{V}^{0:{l_{\max}}}\leftarrow\textrm{Linear}\left(\mathbf{V}^{0:{l_{\max}}}\oplus\hat{\mathbf{V}}^{0:{l_{\max}}}\right)$ $\mathbf{P}\leftarrow\mathbf{P}_{i}+10^{-3}\;\textrm{Linear}\left(\mathbf{V}^{l=1}\right)$ return $(\mathbf{P},\mathbf{V}^{0:{l_{\max}}})$
@@ 3.5 Deep Coarsening Autoencoder
We compose Space Convolution, Self-Interaction and Sequence Convolution modules to define a coarsening/refining block. The block mixes representations across all relevant axes of the domain while producing coarsened, downsampled positions and mixed embeddings. The reverse result – finer positions and decoupled embeddings – is achieved by changing the standard Sequence Convolution to its transpose counterpart. When employing sequence convolutions of stride $S>1$, we increase the dimensionality of the feature representation according to a rescaling factor hyperparameter (Table 1). We stack $L$ coarsening blocks to build a deep neural encoder $\mathcal{E}$, and symetrically $L$ refining blocks to build a decoder $\mathcal{D}$.
3.6 Reconstruction Losses
@@
Vector Map Loss. We train the model by directly comparing internal three-dimensional vector difference maps. Let $V(\mathbf{P})$ denote the internal vector map between all atoms $\mathbf{P}$ in our data, that is, $V(\mathbf{P})^{i,j}=(\mathbf{P}^{i}-\mathbf{P}^{j})\in\mathbb{R}^{N\times N\times 3}$. We then directly minimize $\textrm{HuberLoss}(V(\mathbf{P}),V(\hat{\mathbf{P}}))$ [Huber (1992)].
Cross Entropy Loss. We train the model to predict logits over alphabet $\mathcal{R}$ per residue. We then minimize the cross entropy between predicted logits and ground labels.
Chemical Losses. We incorporate $L_{2}$-norm losses for comparing bonds, angles and dihedrals between prediction and ground truth. For nonbonded atoms, a clash loss is evaluated using standard Van der Waals atomic radii.
An additional stage is employed for processing permutation symmetry breaks. Please refer to Appendix D for further details.
@@ 4.1 Architecture Ablation
Table 1: Autoencoder Recovery from Compressed Representations
@@ 4.2 Latent Conformational Interpolation
Figure 3: Latent Conformational Interpolation. (a) Comparison of chemical validity metrics across interpolated structures of Nuclear Factor of Activated T-Cell (NFAT) Protein (PDB ID: 1S9K and PDB ID: 2O93). We plot results for interpolations on the original data space, on the latent space, and on the autoencoder-reconstructed data space. We note that near interpolation endpoints, latent and AE-reconstructed interpolations perform worse than direct interpolations. This is expected and is a result of the low resolution decoding of the AE. However, only for the latent interpolation the chemical deviation stays consistent, demonstrating that our latent space compactly and smoothly represents the protein space. (b) Linear interpolation of coordinates between two conformational endpoints. (c) Decoded linear latent interpolation between the same encoded states as in (b).
@@
To demonstrate the power of Ophiuchus’s geometric latent space, we show smooth interpolation between two states of a protein structure without explicit latent regularization [Ramaswamy et al. (2021)]. We utilize PDBFlex dataset [Hrabe et al. (2016)] to pick pairs of flexible conformational snapshots of a protein. We utilize these pairs as endpoints for our interpolation. We train a large Ophiuchus reconstruction model on general PDB data. The model coarsens up to 485-residue proteins into a single high-order geometric representation using 6 convolutional downsampling layers each with kernel size 5 and stride 3. The endpoint structures are compressed into single high-order geometric representation to enable direct latent interpolation. We then compare the results of linear interpolation in this latent space against linear interpolation in the coordinate domain (Fig. 3). To determine chemical validity of intermediate states, we scan protein data to quantify average bond lengths and inter-bond angles. We then calculate $L_{2}$ deviation of bonds and angles of interpolated structures. Additionally, we measure atomic clashes by counting collisions of van der Waals radii of non-bonded atoms. Our experiment shows that the decoded structures from latent interpolation maintain a consistent profile of chemical validity, while direct interpolation in the coordinate domain disrupts it significantly. We provide an additional example in Appendix G.
[…]
4.3 Latent Diffusion
We investigate the performance of a generative model trained on the Ophiuchus latent space for sampling decodable protein structure representations. We leverage the miniprotein dataset produced by [Cao et al. (2022)] for protein binder design. The dataset consists of 66k all-atom structures of sequence length between 50 and 65, and composes of diverse folds and sequences across 5 secondary structure topologies. The small sequence sizes facilitate precise reconstruction from unified bottlenecks, enabling a generation method exclusively on latent feature representations to implicitly generate all-atom coordinates.
To that end, we employ DDPMs [Ho et al. (2020); Sohl-Dickstein et al. (2015)] on signals over SO(3) to produce decodable embeddings. First, we train deep autoencoder models for backbone and all-atom structures over the miniprotein dataset. In order to ensure that bottleneck representations $\mathbf{Z}_{0}$ are well-behaved targets for generation, we further regularize the latent space. We build a denoising network by stacking $L_{D}=7$ layers of the Ophiuchus Self-Interaction Fig. (4.a). We parameterize the diffusion model following [Salimans & Ho (2022)], and train it to recover $\mathbf{Z}_{0}$ from $\mathbf{Z}_{t}$: $\mathcal{L}_{\textrm{diffusion}}=\mathbb{E}_{{\epsilon},t}\big{[}w(\lambda_{t})||\hat{\mathbf{Z}}_{0}(\mathbf{Z}_{t})-\mathbf{Z}_{0}||^{2}_{2}\big{]}$. (Appendix H)
In Fig.(5.b) we show the distribution of TM-scores for joint sampling of sequence and all-atom structure by the diffusion model. We produce marginal distributions of generated samples Fig.(5.e) and find them to successfully approximate the densities of ground truth data. To test the robustness of joint sampling of structure and sequence, we compute self-consistency TM-Scores[Trippe et al. (2023)]. 54% of our sampled backbones have scTM scores > 0.5 compared to 77.6% of samples from RFDiffusion. We also sample a protein between 50-65 amino acids in an average 0.21s compared to 11s taken by RFDiffusion on a RTX A6000. Details can be found in Appendix H.
@@ 5 Conclusion and Future Work
In this work, we introduced a new unsupervised autoencoder model for protein structure and sequence representation. Through extensive ablation on its architecture, we quantified the trade-offs between model complexity and representation quality. We demonstrated the power of our learned representations in latent interpolation, and investigated its usage as basis for efficient latent generation of all-atom proteins. Our studies suggest Ophiuchus to provide a strong foundation for constructing state-of-the-art protein neural architectures. In future work, we will investigate scaling Ophiuchus representations and generation to larger proteins and additional molecular domains.

### v2

@@ 3.2 Self-Interaction
Our Self-Interaction is designed to model the internal interactions of atoms within each residue. This transformation updates the feature vectors $\mathbf{V}^{0:{l_{\max}}}_{i}$ centered at the same residue $i$. Importantly, it blends feature vectors $\mathbf{V}^{l}$ of varying degrees $l$ by employing tensor products of the features with themselves. We offer two implementations of these tensor products to cater to different computational needs. Our Self-Interaction module draws inspiration from MACE [Batatia et al. (2022)]. For a comprehensive explanation, please refer to Appendix A.3.
@@
$\mathbf{V}^{0:{l_{\max}}}\leftarrow\mathbf{V}^{0:{l_{\max}}}\oplus\left(\mathbf{V}^{0:{l_{\max}}}\right)^{\otimes 2}$  normal-$\triangleright$ Tensor Square and Concatenate
$\mathbf{V}^{0:{l_{\max}}}\leftarrow\mathrm{Linear}(\mathbf{V}^{0:{l_{\max}}})$  normal-$\triangleright$ Update features
$\mathbf{V}^{0:{l_{\max}}}\leftarrow\textrm{MLP}(\mathbf{V}^{l=0})\cdot\mathbf{V}^{0:{l_{\max}}}$  normal-$\triangleright$ Gate Activation Function
@@ 3.3 Sequence Convolution
To take advantage of the sequential nature of proteins, we propose a one-dimensional, roto-translational equivariant convolutional layer for acting on geometric features and positions of sequence neighbors. Given a kernel window size $K$ and stride $S$, we concatenate representations $\mathbf{V}^{0:{l_{\max}}}_{i-\frac{K}{2}:i+\frac{K}{2}}$ with the same $l$ value. Additionally, we include normalized relative vectors between anchoring positions $\mathbf{P}_{i-\frac{K}{2}:i+\frac{K}{2}}$. Following conventional CNN architectures, this concatenated representation undergoes a linear transformation. The scalars in the resulting representation are then converted into weights, which are used to combine window coordinates into a new coordinate. To ensure translational equivariance, these weights are constrained to sum to one.
@@
$\mathbf{P}\leftarrow\sum_{k=1}^{K}w_{k}\mathbf{P}_{k}$   normal-$\triangleright$ Coarsen coordinates
$\tilde{\mathbf{V}}\leftarrow\bigoplus_{K}\mathbf{V}^{0:{l_{\max}}}_{1:K}$   normal-$\triangleright$ Stack features
$\tilde{\mathbf{P}}\leftarrow\bigoplus^{K,K}_{i=1,j=1}Y\left(\mathbf{P}_{i}-\mathbf{P}_{j}\right)$   normal-$\triangleright$ Stack relative vectors
$\mathbf{V}^{0:{l_{\max}}}\leftarrow\textrm{Linear}\Big{(}\tilde{\mathbf{V}}\oplus\tilde{\mathbf{P}}\Big{)}$   normal-$\triangleright$ Coarsen features and vectors
return $(\mathbf{P},\mathbf{V}^{0:{l_{\max}}})$
@@
When $S>1$, sequence convolutions reduce the dimensionality along the sequence axis, yielding mixed representations and coarse coordinates. To reverse this procedure, we introduce a transpose convolution algorithm that uses its $\mathbf{V}^{l=1}$ features to spawn coordinates. For further details, please refer to Appendix A.6.
@@ 3.4 Spatial Convolution
To capture interactions of residues that are close in three-dimensional space, we introduce the Spatial Convolution. This operation updates representations and positions through message passing within k-nearest spatial neighbors. Message representations incorporate SO(3) signals from the vector difference between neighbor coordinates, and we aggregate messages with a permutation-invariant means. After aggregation, we linearly transform the vector representations into a an update for the coordinates.
@@
$(\tilde{\mathbf{P}},\tilde{\mathbf{V}}^{0:{l_{\max}}})_{1:k}\leftarrow\textrm{$k$-Nearest-Neighbors}(\mathbf{P}_{i},\mathbf{P}_{1:N})$ $R_{1:k},\;\phi_{1:k}\leftarrow\textrm{Embed}(||\tilde{\mathbf{P}}_{1:k}-\mathbf{P}_{i}||_{2}),\;Y(\tilde{\mathbf{P}}_{1:k}-\mathbf{P}_{i})$   normal-$\triangleright$ Edge Features
$\tilde{\mathbf{V}}^{0:{l_{\max}}}_{1:k}\leftarrow\textrm{MLP}(R_{k})\cdot\Big{(}\textrm{Linear}(\tilde{\mathbf{V}}^{0:{l_{\max}}}_{1:k})+\textrm{Linear}\left(\phi_{1:k}\right)\Big{)}$   normal-$\triangleright$ Prepare messages
$\mathbf{V}^{0:{l_{\max}}}\leftarrow\textrm{Linear}\left(\mathbf{V}^{0:{l_{\max}}}_{i}+\frac{1}{k}\left(\sum_{k}\tilde{\mathbf{V}}^{0:{l_{\max}}}_{k}\right)\right)$   normal-$\triangleright$ Aggregate and update
$\mathbf{P}\leftarrow\mathbf{P}_{i}+\textrm{Linear}\left(\mathbf{V}^{l=1}\right)$  normal-$\triangleright$ Update positions
return $(\mathbf{P},\mathbf{V}^{0:{l_{\max}}})$
@@ 3.5 Deep Coarsening Autoencoder
We compose Space Convolution, Self-Interaction and Sequence Convolution modules to define a coarsening/refining block. The block mixes representations across all relevant axes of the domain while producing coarsened, downsampled positions and mixed embeddings. The reverse result – finer positions and decoupled embeddings – is achieved by changing the standard Sequence Convolution to its transpose counterpart. When employing sequence convolutions of stride $S>1$, we increase the dimensionality of the feature representation according to a rescaling factor hyperparameter $\rho$. We stack $L$ coarsening blocks to build a deep neural encoder $\mathcal{E}$ (Alg.7), and symetrically $L$ refining blocks to build a decoder $\mathcal{D}$ (Alg. 8).
3.6 Autoencoder Reconstruction Losses
@@
Vector Map Loss. We train the model by directly comparing internal three-dimensional vector difference maps. Let $V(\mathbf{P})$ denote the internal vector map between all atoms $\mathbf{P}$ in our data, that is, $V(\mathbf{P})^{i,j}=(\mathbf{P}^{i}-\mathbf{P}^{j})\in\mathbb{R}^{3}$. We define the vector map loss as $\mathcal{L}_{\textrm{VectorMap}}=\textrm{HuberLoss}(V(\mathbf{P}),V(\hat{\mathbf{P}}))$ [Huber (1992)]. When computing this loss, an additional stage is employed for processing permutation symmetry breaks. More details can be found in Appendix B.1.
Residue Label Cross Entropy Loss. We train the model to predict logits $\bm{\ell}$ over alphabet $\mathcal{R}$ for each residue. We use the cross entropy between predicted logits and ground labels: $\mathcal{L}_{\textrm{CrossEntropy}}=\textrm{CrossEntropy}(\bm{\ell},\mathbf{R})$
Chemistry Losses. We incorporate $L_{2}$ norm-based losses for comparing bonds, angles and dihedrals between prediction and ground truth. For non-bonded atoms, a clash loss is evaluated using standard Van der Waals atomic radii (Appendix B.2)
Please refer to Appendix B for further details on the loss.
3.7 Latent Diffusion
We train an SO(3)-equivariant DDPM [Ho et al. (2020); Sohl-Dickstein et al. (2015)] on the latent space of our autoencoder. We pre-train an autoencoder and transform each protein $\mathbf{X}$ from the dataset into a geometric tensor of irreducible representations of SO(3): $\mathbf{Z}=\mathcal{E}(\mathbf{X})$. We attach a diffusion process of $T$ steps on the latent variables, making $\mathbf{Z}_{0}=\mathbf{Z}$ and $\mathbf{Z}_{T}=\mathcal{N}(0,1)$. We follow the parameterization described in [Salimans & Ho (2022)], and train a denoising model to reconstruct the original data $\mathbf{Z}_{0}$ from its noised version $\mathbf{Z}_{t}$:
$\mathcal{L}_{\textrm{diffusion}}=\mathbb{E}_{{\epsilon},t}\big{[}w(\lambda_{t})||\hat{\mathbf{Z}}_{0}(\mathbf{Z}_{t})-\mathbf{Z}_{0}||^{2}_{2}\big{]}$
In order to ensure that bottleneck representations $\mathbf{Z}_{0}$ are well-behaved for generation purposes, we regularize the latent space of our autoencoder (Appendix B.3). We build a denoising network with $L_{D}$ layers of Self-Interaction. Please refer to Appendix F for further details.
@@ 4 Methods and Results
4.1 Autoencoder Architecture Comparison
We compare Ophiuchus to the architecture proposed in [Fu et al. (2023)], which uses the EGNN-based architecture [Satorras et al. (2022)] for autoencoding protein backbones. To the best of our knowledge, this is the only other model that attempts protein reconstruction in three-dimensions with roto-translation equivariant networks. For demonstration purposes, we curate a small dataset of protein $\mathbf{C}_{\alpha}$-backbones from the PDB with lengths between 16 and 64 and maximum resolution of up to 1.5 Å, resulting in 1267 proteins. We split the data into train, validation and test sets with ratio [0.8, 0.1, 0.1]. In table 1, we report the test performance at best validation step, while avoiding over-fitting during training.
Table 1: Reconstruction from single feature bottleneck
[…]
We find that Ophiuchus vastly outperforms the EGNN-based architecture. Ophiuchus recovers the protein sequence and $\mathbf{C}_{\alpha}$ backbone with significantly better accuracy, while using orders of magnitude less parameters. Refer to Appendix C for further details.
@@ 4.2 Architecture Ablation
Table 2: Ophiuchus Recovery from Compressed Representations - All-Atom
@@
Figure 3: Protein Latent Diffusion. (a) We attach a diffusion process to Ophiuchus representations and learn a denoising network to sample embeddings of SO(3) that decode to protein structures. (b) Random samples from 485-length backbone model.
@@ 4.3 Latent Conformational Interpolation
To demonstrate the power of Ophiuchus’s geometric latent space, we show smooth interpolation between two states of a protein structure without explicit latent regularization (as opposed to [Ramaswamy et al. (2021)]). We use the PDBFlex dataset [Hrabe et al. (2016)] and pick pairs of flexible proteins. Conformational snapshots of these proteins are used as the endpoints for the interpolation. We train a large Ophiuchus reconstruction model on general PDB data. The model coarsens up to 485-residue proteins into a single high-order geometric representation using 6 convolutional downsampling layers each with kernel size 5 and stride 3. The endpoint structures are compressed into single geometric representation which enables direct latent interpolation in feature .
We compare the results of linear interpolation in the latent space against linear interpolation in the coordinate domain (Fig. 11). To determine chemical validity of intermediate states, we scan protein data to quantify average bond lengths and inter-bond angles. We calculate the $L_{2}$ deviation from these averages for bonds and angles of interpolated structures. Additionally, we measure atomic clashes by counting collisions of Van der-Waals radii of non-bonded atoms (Fig. 11). Although the latent and autoencoder-reconstructed interpolations perform worse than direct interpolation near the original data points, we find that only the latent interpolation structures maintain a consistent profile of chemical validity throughout the trajectory, while direct interpolation in the coordinate domain disrupts it significantly. This demonstrates that the learned latent space compactly and smoothly represents protein conformations.
4.4 Latent Diffusion Experiments and Benchmarks
Our ablation study (Tables 2 and 4) shows successful recovery of backbone structure of large proteins even for large coarsening rates. However, we find that for sequence reconstruction, larger models and longer training times are required. During inference, all-atom models rely on the decoding of the sequence, thus making significantly harder for models to resolve all-atom generation. Due to computational constraints, we investigate all-atom latent diffusion models for short sequence lengths, and focus on backbone models for large proteins. We train the all-atom models with mini-proteins of sequences shorter than 64 residues, leveraging the MiniProtein scaffolds dataset produced by [Cao et al. (2022)]. In this regime, our model is precise and successfully reconstructs sequence and all-atom positions. We also instantiate an Ophiuchus model for generating the backbone trace for large proteins of length 485. For that, we train our model on the PDB data curated by [Yim et al. (2023)]. We compare the quality of diffusion samples from our model to RFDiffusion (T=50) and FrameDiff (N=500 and noise=0.1) samples of similar lengths. We generated 500 unconditional samples from each of the models for evaluation.
In Table 3 we compare sampling the latent space of Ophiuchus to existing models. We find that our model performs comparably in terms of different generated sample metrics, while enabling orders of magnitude faster sampling for proteins. For all comparisons we run all models on a single RTX6000 GPU. Please refer to Appendix F for more details.
Table 3: Comparison to different diffusion models.
Model Dataset Sampling Time (s) $\downarrow$ scRMSD (< 2Å) $\uparrow$ scTM (>0.5) $\uparrow$ Diversity $\uparrow$ FrameDiff [Yim et al. (2023)] PDB 8.6 0.17 0.81 0.42 RFDiffusion [Trippe et al. (2023)] PDB + AlphaFold DB 50 0.79 0.99 0.64 Ophiuchus-64 All-Atom MiniProtein 0.15 0.32 0.56 0.72 Ophiuchus-485 Backbone PDB 0.46 0.18 0.36 0.39
@@ 5 Conclusion and Future Work
In this work, we introduced a new autoencoder model for protein structure and sequence representation. Through extensive ablation on its architecture, we quantified the trade-offs between model complexity and representation quality. We demonstrated the power of our learned representations in latent interpolation, and investigated its usage as basis for efficient latent generation of backbone and all-atom protein structures. Our studies suggest Ophiuchus to provide a strong foundation for constructing state-of-the-art protein neural architectures. In future work, we will investigate scaling Ophiuchus representations and generation to larger proteins and additional molecular domains.